Can someone please help me with this I have to write this fraction in simplest form
-15x^3 y^3/-20xy^4

Answer by checkley71(8403)

-15X^3Y^3/-20XY^4
FACTORING THE INTEGERS WE GET BY DIVIDING BOTH BY -5 (-15/-20=3/4)
factoring the x terms x^3/x we get x^2 x^[3-1]=x^2
factoring the y terms y^3/y^4 we get 1/y y^[3-4)=y^-1 or 1/y
3X^2/4Y ANSWER.